OpenStudy (anonymous):

What is the value of 54 – 4a2 + 3b3 when a = −2 and b = 4?

OpenStudy (anonymous):

84

OpenStudy (k_lynn):

You will need to add in the variables first...now your equation will look like this > 54-4*(-2)*2+3*4*3. I added the * in there so you will know it means to multiply. Remember PEMDAS and do the multiplication first. 4*(-2)=-8. -8*2=-16. Remember that number because you will need it later. 3*4=12. 12*3=36. Remember that number also. Now you can do the equation left to right like this > 54-(-16)+36. 54-(-16)=70. 70+36=106. Therefore, the answer to this equation is 106. I hope this helped you understand!